. What is the relation between nanosecond and second

Grade:7

5 Answers

Arun
6 years ago
A nanosecond is one billionth of a second. Seconds. A second is the base unit of time. Originally, it was based on the length of the day.1 sec = 10^9 nano seconds
Francis Reddy
5 years ago
one thousand-millionth of a second.
A nanosecond is an SI unit of time equal to one thousand-millionth of a second, that is, ¹/1 000 000 000 of a second, or 10−9 seconds. The term combines the prefix nano- with the basic unit for one-sixtieth of a minute.
 
Scholr Boy
5 years ago
Use Scholr App for solving doubts
A nanosecond (ns) is an SI unit of time equal to one thousand-millionth of a second (or one billionth of a second), that is, 1/1,000,000,000 of a second, or 109 second
Sayantan Garai
5 years ago
1ns= 10-9s
ns--nanosecond
s--second
n=10-9 is an S.I. prefix.
s-- unit of time.
 
 
 
__________________________________
 
 
 
Ayan
5 years ago
A nanosecond is one billionth of a second. Seconds.1 sec = 10^9 nano seconds................................................